# Rational Tuning of LLM Cascades via Probabilistic Modeling

### v1

@@ 4.3.2 Testing the Gumbel Copulas for Modeling LLM Correlations
Table 4: Shows the goodness-of-fit of our Gumbel copula models for modeling pairwise correlations between LLMs, based on a Cramér-von Mises ($\sqrt{n}\text{CvM}$) test using parametric bootstrapping. We now report the $\sqrt{n}\text{CvM}$ value, the number of null hypothesis rejections (out of 10 and 6 model pairs for the Llama and Qwen & OpenAI groups, respectively), the percentage of rejections, as well as the geometric and arithmetic mean of $p$ values.
@@ 4.4 Rational Tuning of Confidence Thresholds
Grid search baseline: this baseline selects optimal confidence thresholds by searching over adaptive grids computed from the model-specific quantiles of calibrated confidence. Specifically, in each dimension the grid ranges from $\phi_{\text{min}}$ to $\phi_{\text{max}}$ in increments of $2.5\%$ probability mass. After scoring all candidate threshold combinations, we use the Skyline operator (implemented in the Python package paretoset [^5] Open-source implementation available at github.com/tommyod/paretoset. Accessed January 13, 2025.) to filter the candidate threshold vectors down to the Pareto-optimal set (Börzsönyi et al., (2001)). A candidate threshold vector $\boldsymbol{\theta}=(\phi_{1},...,\phi_{k-1})$ is Pareto-optimal if its performance metrics $(\mathbb{P}_{\boldsymbol{\theta}}(\text{Correct}),\mathbb{E}_{\boldsymbol{\theta}}[\text{Cost}])$ are not dominated by any other candidate threshold vector $\boldsymbol{\theta^{\prime}}$ in the sense that $\mathbb{P}_{\boldsymbol{\theta^{\prime}}}(\text{Correct})>\mathbb{P}_{\boldsymbol{\theta}}(\text{Correct})$ and $\mathbb{E}_{\boldsymbol{\theta^{\prime}}}[\text{Cost}]<\mathbb{E}_{\boldsymbol{\theta}}[\text{Cost}]$.
@@ 5 Conclusion
Proposition.
2. Let $C$ be a cascade $M_{1}\rightarrow...\rightarrow M_{k}$ with confidence thresholds $(\phi_{1},...,\phi_{k-1})$. Assume that the distribution functions for the calibrated confidences $\Phi_{i}$ satisfy (5), for $i=1,2,...,k$. Assume further that the expected numbers of input and output tokens, $T_{i}^{\text{(in)}}$ and $T_{i}^{\text{(out)}}$, for each model $i$ are independent of the calibrated confidences $\Phi_{1},...,\Phi_{k}$. Then the probability of correctness $\mathbb{P}(\text{correct})$ and expected cost $\mathbb{E}[Cost]$ for the cascade $C$ are
$\mathbb{P}(\text{correct})=$ $\sum_{i=1}^{k}\mathbb{P}(\Phi_{1}\leq\phi_{1})\left(\prod_{j=2}^{i-1}\mathbb{P}(\Phi_{j}\leq\phi_{j}|\Phi_{j-1}\leq\phi_{j-1})\right)\int_{\{\Phi_{i}>\phi_{i}\}}\Phi_{i}(\omega)\text{~{}d}\mathbb{P}(\omega|\Phi_{i-1}\leq\phi_{i-1})$ (21)
$\mathbb{E}[Cost]=$ $\sum_{i=1}^{k}\mathbb{P}(\Phi_{1}\leq\phi_{1})\left(\prod_{j=2}^{i-1}\mathbb{P}(\Phi_{j}\leq\phi_{j}|\Phi_{j-1}\leq\phi_{j-1})\right)(1-\mathbb{P}(\Phi_{i}\leq\phi_{i}|\Phi_{i-1}\leq\phi_{i-1}))\sum_{j=1}^{i}\mathbb{E}[C_{j}],$ (22)
where $C_{i}$ is the cost per query of model $i$. Specifically, if $\gamma_{i}^{(\text{in})}$ and $\gamma_{i}^{(\text{out})}$ are the costs per input and output token, $C_{i}=\gamma_{i}^{(\text{in})}T_{i}^{\text{(in)}}+\gamma_{i}^{(\text{out})}T_{i}^{\text{(out)}}$.
@@
Hence, the problem reduces to computing $\mathbb{E}[\Phi_{i}\mathds{1}\{\tau=i\}]$ for each model $i$. This is the integral of $\Phi_{i}$ over the set $\{\tau=i\}$. We have
$\mathbb{E}[\Phi_{i}\mathds{1}\{\tau=i\}]=$ $\int\Phi_{i}\mathds{1}_{\{\Phi>\phi_{i}\}}\prod_{j=2}^{i-1}\mathds{1}_{\{\Phi_{j}\leq\phi_{j}\}}\text{~{}d}\mathbb{P}$ (24)
$=$ $~{}\mathbb{P}(\Phi_{1}\leq\phi_{1},...,\Phi_{i-1}\leq\phi_{i-1})\int\Phi_{i}\mathds{1}_{\{\Phi>\phi_{i}\}}\frac{\prod_{j=2}^{i-1}\mathds{1}_{\{\Phi_{j}\leq\phi_{j}\}}}{\mathbb{P}(\Phi_{1}\leq\phi_{1},...,\Phi_{i-1}\leq\phi_{i-1})}\text{~{}d}\mathbb{P}$ (25)
$=$ $~{}\mathbb{P}(\Phi_{1}\leq\phi_{1})\prod_{j=2}^{i-1}\mathbb{P}(\Phi_{j}\leq\phi_{j}|\Phi_{j-1}\leq\phi_{j-1})\int_{\{\Phi>\phi_{i}\}}\Phi_{i}\text{~{}d}\mathbb{P}(\cdot|\cap_{j=1}^{i-1}\{\Phi_{j}\leq\phi_{j}\})$ (26)
$=$ $~{}\mathbb{P}(\Phi_{1}\leq\phi_{1})\prod_{j=2}^{i-1}\mathbb{P}(\Phi_{j}\leq\phi_{j}|\Phi_{j-1}\leq\phi_{j-1})\int_{\{\Phi>\phi_{i}\}}\Phi_{i}\text{~{}d}\mathbb{P}(\cdot|\Phi_{i-1}\leq\phi_{i-1}),$ (27)
where in the last line we used the Markov assumption (5). This concludes the proof of the formula for the probability of correctness. To obtain the formula for the expected cost, note that the integral
@@
0:  Confidence Thresholds $(\phi_{1},...,\phi_{k})$
[…]
2:  $\text{D}\leftarrow 1$  # cumulative transition probability
@@
5:
6:  A.append($\mathbb{P}(\Phi_{1}>\phi_{1})$)
7:  C.append($\text{Cost}\times\mathbb{P}(\Phi_{1}>\phi_{1})$)
8:  $\text{D}\leftarrow\text{D}\times\mathbb{P}(\Phi_{1}\leq\phi_{1})$
9:
10:  for $i=2$$k$ do
11:     $\text{Cost}\leftarrow\text{Cost}+\mathbb{E}[C_{i}]$
12:     A.append($\text{D}\times\int_{\{\Phi_{i}>\phi_{i}\}}\Phi_{i}(\omega)\text{~{}d}\mathbb{P}(\omega|\Phi_{i-1}\leq\phi_{i-1})$)
13:     C.append($\text{D}\times(1-\mathbb{P}(\Phi_{i}\leq\phi_{i}|\Phi_{i-1}\leq\phi_{i-1}))\times\text{Cost}$)
14:     $\text{D}\leftarrow\text{D}\times\mathbb{P}(\Phi_{i}\leq\phi_{i}|\Phi_{i-1}\leq\phi_{i-1})$
15:  end for
16:
17:  $\mathbb{P}(\text{Correct})\leftarrow\text{sum}(\text{A})$
18:  $\mathbb{E}[\text{Cost}]\leftarrow\text{sum}(\text{C})$
19:
20:  return  ($\mathbb{P}(\text{Correct}),\mathbb{E}[\text{Cost}]$)

### v3

@@ 4.3.2 Testing the Gumbel Copulas for Modeling LLM Correlations
Table 4: Shows the goodness-of-fit of our Gumbel copula models for modeling pairwise correlations between LLMs, based on a Cramér-von Mises ($\sqrt{n}\text{CvM}$) test using parametric bootstrapping. We report the $\sqrt{n}\text{CvM}$ value, the number of null hypothesis rejections (out of 10 and 6 model pairs for the Llama and Qwen & OpenAI groups, respectively), the percentage of rejections, as well as the geometric and arithmetic mean of $p$ values.
@@ 4.4 Rational Tuning of Confidence Thresholds
Grid search baseline: this baseline selects optimal confidence thresholds by searching over adaptive grids computed from the model-specific quantiles of calibrated confidence. Specifically, in each dimension the grid ranges from $\phi_{\text{min}}$ to $\phi_{\text{max}}$ in increments of $2.5\%$ probability mass. After scoring all candidate threshold combinations, we use the Skyline operator (implemented in the Python package paretoset [^5] Open-source implementation available at github.com/tommyod/paretoset. Accessed January 13, 2025.) to filter the candidate threshold vectors down to the Pareto-optimal set (Börzsönyi et al.,, 2001). A candidate threshold vector $\boldsymbol{\theta}=(\phi_{1},...,\phi_{k-1})$ is Pareto-optimal if its performance metrics $(\mathbb{P}_{\boldsymbol{\theta}}(\text{Correct}),\mathbb{E}_{\boldsymbol{\theta}}[\text{Cost}])$ are not dominated by any other candidate threshold vector $\boldsymbol{\theta^{\prime}}$ in the sense that $\mathbb{P}_{\boldsymbol{\theta^{\prime}}}(\text{Correct})>\mathbb{P}_{\boldsymbol{\theta}}(\text{Correct})$ and $\mathbb{E}_{\boldsymbol{\theta^{\prime}}}[\text{Cost}]<\mathbb{E}_{\boldsymbol{\theta}}[\text{Cost}]$.
@@ 5 Conclusion
Proposition 2.
Consider a cascade $M_{1}\rightarrow...\rightarrow M_{k}$ with confidence thresholds $(\phi_{1},...,\phi_{k-1})$. Assume that the distribution functions for the calibrated confidences $\Phi_{i}$ satisfy (5), for $i=1,2,...,k$. Assume further that the expected numbers of input and output tokens, $T_{i}^{\text{(in)}}$ and $T_{i}^{\text{(out)}}$, for each model $i$ are independent of the calibrated confidences $\Phi_{1},...,\Phi_{k}$. Then the probability of correctness $\mathbb{P}(\text{Correct})$ and expected cost $\mathbb{E}[Cost]$ for the cascade are
$\mathbb{P}(\text{Correct})=$$\int_{\{\Phi_{1}>\phi_{1}\}}\Phi_{1}(\omega)\text{~{}d}\mathbb{P}(\omega)$ (21)
$+\sum_{i=2}^{k}\mathbb{P}(\Phi_{1}\leq\phi_{1})\left(\prod_{j=2}^{i-1}\mathbb{P}(\Phi_{j}\leq\phi_{j}|\Phi_{j-1}\leq\phi_{j-1})\right)\int_{\{\Phi_{i}>\phi_{i}\}}\Phi_{i}(\omega)\text{~{}d}\mathbb{P}(\omega|\Phi_{i-1}\leq\phi_{i-1})$
$\mathbb{E}[Cost]=$ $~{}(1-\mathbb{P}(\Phi_{1}\leq\phi_{1}))~{}\mathbb{E}[C_{1}]$ (22)
$+\sum_{i=2}^{k}\mathbb{P}(\Phi_{1}\leq\phi_{1})\left(\prod_{j=2}^{i-1}\mathbb{P}(\Phi_{j}\leq\phi_{j}|\Phi_{j-1}\leq\phi_{j-1})\right)(1-\mathbb{P}(\Phi_{i}\leq\phi_{i}|\Phi_{i-1}\leq\phi_{i-1}))\sum_{j=1}^{i}\mathbb{E}[C_{j}],$
where $C_{i}$ is the cost per query of model $i$. Specifically, if $\gamma_{i}^{(\text{in})}$ and $\gamma_{i}^{(\text{out})}$ are the costs per input and output token, $C_{i}=\gamma_{i}^{(\text{in})}T_{i}^{\text{(in)}}+\gamma_{i}^{(\text{out})}T_{i}^{\text{(out)}}$. To simplify the notation, we let $\phi_{k}:=-\infty$ (although there is no confidence threshold for the final model in the cascade).
@@
Hence, the problem reduces to computing $\mathbb{E}[\Phi_{i}\mathds{1}\{\tau=i\}]$ for each model $i$. This is the integral of $\Phi_{i}$ over the set $\{\tau=i\}$. For $i\geq 2$, we have
$\mathbb{E}[\Phi_{i}\mathds{1}\{\tau=i\}]=$ $\int\Phi_{i}\mathds{1}_{\{\Phi_{i}>\phi_{i}\}}\prod_{j=1}^{i-1}\mathds{1}_{\{\Phi_{j}\leq\phi_{j}\}}\text{~{}d}\mathbb{P}$ (24)
$=$ $~{}\mathbb{P}(\Phi_{1}\leq\phi_{1},...,\Phi_{i-1}\leq\phi_{i-1})\int\Phi_{i}\mathds{1}_{\{\Phi_{i}>\phi_{i}\}}\frac{\prod_{j=1}^{i-1}\mathds{1}_{\{\Phi_{j}\leq\phi_{j}\}}}{\mathbb{P}(\Phi_{1}\leq\phi_{1},...,\Phi_{i-1}\leq\phi_{i-1})}\text{~{}d}\mathbb{P}$ (25)
$=$ $~{}\mathbb{P}(\Phi_{1}\leq\phi_{1})\prod_{j=2}^{i-1}\mathbb{P}(\Phi_{j}\leq\phi_{j}|\Phi_{j-1}\leq\phi_{j-1})\int_{\{\Phi_{i}>\phi_{i}\}}\Phi_{i}\text{~{}d}\mathbb{P}(\cdot|\cap_{j=1}^{i-1}\{\Phi_{j}\leq\phi_{j}\})$ (26)
$=$ $~{}\mathbb{P}(\Phi_{1}\leq\phi_{1})\prod_{j=2}^{i-1}\mathbb{P}(\Phi_{j}\leq\phi_{j}|\Phi_{j-1}\leq\phi_{j-1})\int_{\{\Phi_{i}>\phi_{i}\}}\Phi_{i}\text{~{}d}\mathbb{P}(\cdot|\Phi_{i-1}\leq\phi_{i-1}).$ (27)
To obtain Equation (26), we applied the Markov assumption (5) and switched from the standard probability measure $\mathbb{P}(\cdot)$ to the conditional probability measure $\mathbb{P}(\cdot\cap A)/\mathbb{P}(A)$, where $A=\{\Phi_{1}\leq\phi_{1},...,\Phi_{i-1}\leq\phi_{i-1}\}$. To obtain the last line, we applied the Markov assumption (5) again.
For $i=1$, we have that
$\mathbb{E}[\Phi_{1}\mathds{1}\{\tau=1\}]=\int_{\{\Phi_{1}>\phi_{1}\}}\Phi_{1}(\omega)\text{~{}d}\mathbb{P}(\omega).$ (28)
This concludes the proof of the formula for the probability of correctness. To obtain the formula for the expected cost, we reason analogously and note that the integral
@@
0:  confidence thresholds $\phi_{1},...,\phi_{k-1}\in\mathbb{R}^{k-1}$
@@
[…]
5:  $\phi_{k}\leftarrow-\infty$
6:
7:  correctness_terms.append($\int_{\{\Phi_{1}>\phi_{1}\}}\Phi_{1}(\omega)\texttt{~{}d}\mathbb{P}(\omega)$)
[…]
10:
11:  for $i=2$$k$ do
[…]
16:  end for
17:
[…]
20:
21:  return  ($\mathbb{P}(\text{Correct}),\mathbb{E}[\text{Cost}]$)